OpenStudy (anonymous):

What is the common ratio of the geometric sequence: 1/8, -1/3, 2, -12...

jimthompson5910 (jim_thompson5910):

Take the second term -1/3 and divide it by the first term 1/8 what do you get? \[\Large -\frac{1}{3} \div \frac{1}{8} = ???\]

OpenStudy (asnaseer):

Your sequence does not look quite right. It would make more sense if the first term was \(1/18\) instead of \(1/8\)

OpenStudy (anonymous):

You would get: -11/24

jimthompson5910 (jim_thompson5910):

Is the first term 1/18 instead of 1/8 ?

OpenStudy (anonymous):

No, the question has it 1/18.

jimthompson5910 (jim_thompson5910):

so the sequence is really this 1/18, -1/3, 2, -12... right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok so compute the following below \[\Large -\frac{1}{3} \div \frac{1}{18} = ???\]

jimthompson5910 (jim_thompson5910):

flip the second fraction and then multiply

OpenStudy (anonymous):

Okay. Is this the last step? If so, then I can do the rest.

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

-18/3

jimthompson5910 (jim_thompson5910):

which reduces to what?

OpenStudy (anonymous):

-6/1

jimthompson5910 (jim_thompson5910):

which is just -6

jimthompson5910 (jim_thompson5910):

now do the same for the third term and second term \[\Large 2 \div -\frac{1}{3} = ???\]

OpenStudy (anonymous):

-2/3

jimthompson5910 (jim_thompson5910):

no, again first flip the second fraction and multiply

jimthompson5910 (jim_thompson5910):

that's the rule for dividing fractions

jimthompson5910 (jim_thompson5910):

think of 2 as 2/1

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Thanks for the help. I've figured it out.

jimthompson5910 (jim_thompson5910):

you should keep getting -6 each time you divide any term by its previous term